# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

### v2

@@ 6 Conclusions
In existing MoE language models, the widely-used dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially cause the model to encounter underfitting during training, especially for infrequent tokens. Although fixed routing like Hash Layer can mitigate the routing fluctuation issue, it struggles with capturing rich features for frequent tokens. To tackle the routine fluctuation issue and the representation diversity learning issue at the meantime, we propose a routing masking method termed MaskMoE, aiming at ensuring adequate learning of infrequent tokens while maintaining rich representations learning for frequent tokens. Extensive experiments demonstrate that the proposed MaskMoE outperforms existing MoE methods like SMoE, Hash Layer, and Share-MoE on various downstream tasks and significantly reduces PPL on the Pile validation sets. That will demonstrates the effectiveness of MaskMoE.

### v4

@@ 6 Conclusions
In existing MoE language models, the widely-used dynamic routing methods such as SMoE, exhibit routing fluctuations, and route the occurrences of a token to multiple experts during training, potentially causing the model to encounter underfitting, especially for infrequent tokens. Although fixed routing like Hash Layer can mitigate the routing fluctuation issue, it struggles with capturing rich features for frequent tokens. Considering both routing fluctuation and representation diversity, we propose MaskMoE, a routing masking method that ensures adequate learning for infrequent tokens and rich representations for frequent tokens. Extensive experiments show that MaskMoE outperforms SMoE, Hash Layer, and Share-MoE on various downstream tasks and significantly reduces PPL on the Pile validation sets.